# Large Language Models Are Cross-Lingual Knowledge-Free Reasoners

### v2

@@ Abstract
 [^1] Our code and data is available at: https://github.com/NJUNLP/Knowledge-Free-Reasoning.
@@ 1 Introduction
In this study, we divide a general reasoning task into two separated components: knowledge retrieval and knowledge-free reasoning. The former means recalling certain information from pre-training, while the latter refers to organizing the given knowledge in the context to perform inference and derive a final answer. Figure 1 provides a clearer understanding of these two components and illustrates the cross-lingual transfer issues explored in this paper.
@@ 2.1 Overview
We also aim to investigate the cross-lingual transfer of knowledge-free reasoning, which is less explored in previous work. However, existing reasoning datasets often contain some degree of knowledge retrieval. For instance, while StrategyQA provides knowledge required for reasoning, it is not always complete. Similarly, certain mathematical datasets, like ASDiv, require knowledge retrieval for some problems (as demonstrated in Appendix G). This introduces noise when evaluating the cross-lingual transfer of knowledge-free reasoning. To address this, we constructed a new dataset, the Knowledge-Free Reasoning Dataset (KFRD), which entirely eliminates the need for knowledge retrieval. In addition, we selected several existing datasets that, to the best extent possible, meet the requirements of knowledge-free reasoning to further validate our conclusions. A more detailed explanation of why we constructed KFRD and the dataset selection criteria can be found in Appendix G.
@@ 2.2.1 Reasoning dataset with variable knowledge retrieval demand
We adapt the popular commonsense reasoning datasets, StrategyQA (Geva et al., 2021) and QASC (Khot et al., 2020), to analyze the impact of knowledge retrieval on cross-lingual transfer. They provide pieces of evidence from Wikipedia for answering the question. Examples can be found in Table A5.
@@ 4.1 Impact of knowledge retrieval demand on cross-lingual transfer
We analyze the impact of the amount of knowledge retrieve on cross-lingual transfer in different settings of the reasoning dataset. The results of StrategyQA for the cross-lingual transfer ratio are presented in Figure 2, while the accuracy results are detailed in Figure A2.
@@ 4.2 The cross-lingual transfer of knowledge-free reasoning
For some low-resource languages, such as Hebrew and Arabic in LLaMA-2, German and Hebrew in BLOOMZ  [^2] For the language distribution, please see Appendix F., the cross-lingual transferability is significantly poorer. We hypothesize that this may be due to the model’s weaker language proficiency in these languages, which negatively impacts its transferability. Further analysis of this issue is provided in the following section.
@@
We manually check the data samples and find that there are some interfering cases that can affect cross-lingual transfer, such as minor knowledge retrieves, translation issues, and counterfactual information, as detailed discussed in the Appendix G.
@@ 4.3.1 Training language proficiency
To evaluate the impact of training language proficiency, based on the data distribution of LLaMA-2 (see Appendix F) and previous experiments, we selected German and Chinese as representatives of high-resource languages, and Arabic and Hebrew as representatives of low-resource languages for training. Then, we train models on the KFRD in these languages and evaluated their performance across the 10 languages. The resulting cross-lingual transfer ratios are presented in Figure 6, while the accuracy results are shown in Figure A6.
@@ 5.1 Overview
Built on previous research (Hu et al., 2024; Gao et al., 2024) and our experiments, we observed that the cross-lingual transferability of knowledge retrieval ability is significantly weaker than that of knowledge-free reasoning. To better understand the reasons behind this difference, we conducted a detailed analysis on model internals using two methods: Cosine Similarity of Hidden States and Neuron Activation. Both of the methods have been widely used to measure text similarity (Reimers and Gurevych, 2019; Malkiel et al., 2022; Wang et al., 2024) and analyze models (Dalvi et al., 2019; Sajjad et al., 2022; Rai and Yao, 2024).
@@ 6 Related Work
Gao et al. (2024) evaluated the cross-lingual transferability of models on multiple reasoning datasets, finding significant variations in transfer performance across different datasets. Furthermore, Hu et al. (2024) found that knowledge transferability remains weak across various settings. Building on their conclusions, we distinguish between the knowledge retrieval and knowledge-free reasoning components and extend the analysis to all reasoning tasks. We include further discussion of related reasoning works in Appendix H.
@@ 7 Conclusion and Discussion
Appendix A Detailed of Dataset
@@ Appendix E Additional Results of Experiment
Appendix F Language Distribution of Model Training Corpora
In this section, we present the language distribution of the pre-training corpora, referencing Table A6 from the LLaMA2 paper (Touvron et al., 2023) and Table A7 from the BLOOM paper (Workshop et al., 2022). Unfortunately, we were unable to locate the corresponding distribution data for Mistral and Qwen.
For LLaMA2, languages such as Arabic and Hebrew were not included in the table, indicating that their proportions are lower than 0.005%, categorizing them as extremely low-resource languages. The other eight languages discussed in the paper are represented. Notably, German and Chinese rank as high-resource languages, accounting for 0.17% and 0.13% of the corpus, respectively, holding the second and fifth highest positions.
For BLOOM, only English, French, Chinese, and Arabic are explicitly listed, while other languages are not reported in the table.
Language Percent Language Percent
en 89.70% uk 0.07%
unknown 8.38% ko 0.06%
de 0.17% ca 0.04%
fr 0.16% sr 0.04%
sv 0.15% id 0.03%
zh 0.13% cs 0.03%
es 0.13% fi 0.03%
ru 0.13% hu 0.03%
nl 0.12% no 0.03%
it 0.11% ro 0.03%
ja 0.10% bg 0.02%
pl 0.09% da 0.02%
pt 0.09% sl 0.01%
vi 0.08% hr 0.01%
Table A6: Language distribution in pretraining data with percentage >= 0.005%. Most data is in English, meaning that Llama 2 will perform best for English-language use cases. The large unknown category is partially made up of programming code data.
Language Size in Bytes Language Size in Bytes
Akan 70,1554 Arabic 74,854,900,600
Assamese 291,522,098 Bambara 391,747
Basque 2,360,470,848 Bengali 18,606,823,104
Catalan 17,792,493,289 Chichewa 1,187,405
chiShona 6,638,639 Chitumbuka 170,360
English 484,953,009,124 Fon 2,478,546
French 208,242,620,434 Gujarati 1,199,986,460
Hindi 24,622,119,985 Igbo 14,078,521
Indonesian 19,972,325,222 isiXhosa 14,304,074
isiZulu 8,511,561 Kannada 2,098,453,560
Kikuyu 359,615 Kinyarwanda 40,428,299
Kirundi 3,272,550 Lingala 1,650,804
Luganda 4,568,367 Malayalam 3,662,571,498
Marathi 1,775,483,122 Nepali 2,551,307,393
Northern Sotho 1,764,506 Odia 1,157,100,133
Portuguese 79,277,543,375 Punjabi 1,572,109,752
Sesotho 751,034 Setswana 1,502,200
Simplified Chinese 261,019,433,892 Spanish 175,098,365,045
Swahili 236,482,543 Tamil 7,989,206,220
Telugu 29,934,071,59 Traditional Chinese 762,489,150
Twi 1,265,041 Urdu 2,781,329,959
Vietnamese 43,709,279,959 Wolof 3,606,973
Xitsonga 707,634 Yoruba 89,695,835
Programming Languages 174,700,245,772
Table A7: Linguistic makeup of the ROOTS corpus. The ROOTS corpus is a multilingual dataset used for pretraining the BLOOM model.
Appendix G Reasons for Creating a New Dataset
@@
Appendix H Additional Works on Cross-lingual Transfer
Devlin (2018) introduced mBERT, advancing cross-lingual transfer by capturing shared linguistic patterns in a unified embedding space, enabling zero-shot transfer without parallel corpora. Similarly, Conneau (2019) showed XLM’s effectiveness in optimizing multilingual embeddings, improving performance in translation and classification tasks. Ansell et al. (2021) proposed composable sparse fine-tuning, selectively fine-tuning sparse parameters across languages to reduce interference and boost performance, especially in low-resource settings, outperforming adapter-based methods in tasks like NER and NLI.
Additionally, Ye et al. (2023) assessed the imbalance of knowledge across different languages in LLMs, observing weak cross-lingual transferability of knowledge. Zhu et al. (2024) discovered that training on translated questions can enhance the cross-lingual transferability of reasoning tasks.

### v3

@@ 1 Introduction
Intuitively, abstract reasoning is considered a language-agnostic ability, and thus it should transfer well across languages. The observed performance differences in general reasoning tasks across languages may arise from other factors. In this study, we divide a general reasoning task into two separated components: knowledge retrieval and knowledge-free reasoning. The former means recalling factual knowledge from pre-training  [^1] Similar to closed-book QA tasks, while the latter refers to organizing the given knowledge in the context to perform inference and derive a final answer [^2] Such as Coin Flip Wei et al. (2022). Figure 1 provides a clearer understanding of these two components and illustrates the cross-lingual transfer issues explored in this paper.
@@ 2.1 Overview
We also aim to investigate the cross-lingual transfer of knowledge-free reasoning, which is less explored in previous work. However, existing reasoning datasets often contain some degree of knowledge retrieval. For instance, while StrategyQA (Geva et al., 2021) provides knowledge required for reasoning, it is not always complete. Similarly, certain mathematical datasets, like ASDiv, require knowledge retrieval for some problems (as demonstrated in Appendix H). This introduces noise when evaluating the cross-lingual transfer of knowledge-free reasoning. To address this, we constructed a new dataset, the Knowledge-Free Reasoning Dataset (KFRD), which entirely eliminates the need for knowledge retrieval. In addition, we selected several existing datasets that, to the best extent possible, meet the requirements of knowledge-free reasoning to further validate our conclusions. A more detailed explanation of why we constructed KFRD and the dataset selection criteria can be found in Appendix H.
@@ 2.2.1 Reasoning dataset with variable knowledge retrieval demand
We adapt the popular commonsense reasoning datasets, StrategyQA and QASC (Khot et al., 2020), to analyze the impact of knowledge retrieval on cross-lingual transfer. They provide pieces of evidence from Wikipedia for answering the question. Examples can be found in Table A5.
@@ 4.1 Impact of knowledge retrieval demand on cross-lingual transfer
We analyze the impact of the amount of knowledge retrieved on cross-lingual transfer in different settings of the reasoning dataset. The results of StrategyQA for the cross-lingual transfer ratio are presented in Figure 2, while the accuracy results are detailed in Figure A2.
@@ 4.2 The cross-lingual transfer of knowledge-free reasoning
For some low-resource languages, such as Hebrew and Arabic in LLaMA-2, German and Hebrew in BLOOMZ  [^3] For the language distribution, please see Appendix G., the cross-lingual transferability is significantly poorer. We hypothesize that this may be due to the model’s weaker language proficiency in these languages, which negatively impacts its transferability. Further analysis of this issue is provided in the following section.
@@
We manually check the data samples and find that there are some interfering cases that can affect cross-lingual transfer, such as minor knowledge retrieves, translation issues, and counterfactual information, as detailed discussed in the Appendix H.
We also evaluate the LLaMA-2-7B-Chat model on MMLU before and after finetuning, in order to address the concerns of over-fitting on the fine-tuned datasets and forgetting the world knowledge, which is detailed in Appendix F.
@@ 4.3.1 Training language proficiency
To evaluate the impact of training language proficiency, based on the data distribution of LLaMA-2 (see Appendix G) and previous experiments, we selected German and Chinese as representatives of high-resource languages, and Arabic and Hebrew as representatives of low-resource languages for training. Then, we train models on the KFRD in these languages and evaluated their performance across the 10 languages. The resulting cross-lingual transfer ratios are presented in Figure 6, while the accuracy results are shown in Figure A6.
@@ 5.1 Overview
Built on previous research (Hu et al., 2024b; Gao et al., 2024) and our experiments, we observed that the cross-lingual transferability of knowledge retrieval ability is significantly weaker than that of knowledge-free reasoning. To better understand the reasons behind this difference, we conducted a detailed analysis on model internals using two methods: Cosine Similarity of Hidden States and Neuron Activation. Both of the methods have been widely used to measure text similarity (Reimers and Gurevych, 2019; Malkiel et al., 2022; Wang et al., 2024) and analyze models (Dalvi et al., 2019; Sajjad et al., 2022; Rai and Yao, 2024).
@@ 6 Related Work
Gao et al. (2024) evaluated the cross-lingual transferability of models on multiple reasoning datasets, finding significant variations in transfer performance across different datasets. Furthermore, Hu et al. (2024a) found that knowledge transferability remains weak across various settings. Building on their conclusions, we distinguish between the knowledge retrieval and knowledge-free reasoning components and extend the analysis to all reasoning tasks. Additionally, Ye et al. (2023) assessed the imbalance of knowledge across different languages in LLMs, observing weak cross-lingual transferability of knowledge. Zhu et al. (2024) discovered that training on translated questions can enhance the cross-lingual transferability of reasoning tasks.
There are also some works focusing on the cross-lingual transfer in the pre-LLM era. Devlin (2018) introduced mBERT, advancing cross-lingual transfer by capturing shared linguistic patterns in a unified embedding space, enabling zero-shot transfer without parallel corpora. Similarly, Conneau (2019) showed XLM’s effectiveness in optimizing multilingual embeddings, improving performance in translation and classification tasks. Ansell et al. (2021) proposed composable sparse fine-tuning, selectively fine-tuning sparse parameters across languages to reduce interference and boost performance, especially in low-resource settings, outperforming adapter-based methods in tasks like NER and NLI.
@@ 7 Conclusion and Discussion
Acknowledgements
We would like to thank the anonymous reviewers for their insightful comments. Shujian Huang is the corresponding author. This work is supported by National Science Foundation of China (No. 62376116, 62176120), research project of Nanjing University-China Mobile Joint Institute, the Fundamental Research Funds for the Central Universities (No. 2024300507).
Appendix A Details of Dataset
@@ Appendix E Additional Results of Experiment
Appendix F Evaluation on MMLU
In order to address the concerns of overfitting on the fine-tuned datasets and forgetting the world knowledge, we also evaluate the LLaMA-2-7B-Chat model on MMLU before and after fine-tuning. The results are listed in Table A6. We find that the model’s MMLU performance shows both increases and decreases after finetuning, with relatively small changes in magnitude. This indicates that we do not cause the model to overfit these datasets and maintain its general capabilities.
Finetuned Dataset MMLU(5-shot)
None 40.88
KFRD Arithmetic 45.53
KFRD Symbolic 43.87
KFRD Logical 45.14
StrategyQA WF-all 46.85
ASDiv 38.31
Coin Flip 43.58
ProofWriter 37.72
Table A6: The accuracy of LLaMA-2-7B-Chat model on MMLU both before and after finetuning
Appendix G Language Distribution of Model Training Corpora
In this section, we discuss the language distribution of the pre-training corpora, from the LLaMA2 paper (Touvron et al., 2023) and the BLOOM paper (Workshop et al., 2022). Unfortunately, we were unable to locate the corresponding distribution data for Mistral and Qwen.
For LLaMA2, languages such as Arabic and Hebrew were not included in the table provided in Touvron et al. (2023) (Table 10), indicating that their proportions are lower than 0.005%, categorizing them as extremely low-resource languages. The other eight languages discussed in the paper are represented. Notably, German and Chinese rank as high-resource languages, accounting for 0.17% and 0.13% of the corpus, respectively, holding the second and fifth highest positions.
For BLOOM, only English, French, Chinese, and Arabic are explicitly listed in the paper (Workshop et al., 2022, Table 1), while other languages are not reported in the paper, indicating they are low-resource languages.
Appendix H Reasons for Creating a New Dataset